OpenStudy (anonymous):

Which is the slope of a line with te coordinates (1,4) and (3,1)?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

thank you

hartnn (hartnn):

your welcome ^_^ if you want to verify your anwer, tell me what u got ...

OpenStudy (anonymous):

i got 2/3

hartnn (hartnn):

it is not entirely correct... numerator = y1-y2 =4-1 =... ?

OpenStudy (anonymous):

so its -2/3

hartnn (hartnn):

no, no ...numerator = 4-1 = 3

OpenStudy (anonymous):

haha im sorry im a little hard to work with and oww yahh i got 3/2

hartnn (hartnn):

1-3 = -2 ist actually -3/2 ...

OpenStudy (anonymous):

thnthank you :)

hartnn (hartnn):

welcome :)